What is the pressure of a mixture of nitrogen (N2) and ocygen (o2) if they partial pressure of N2 is 495 mm Hg and the partial pressure of O2 is 156 mmhg
1
Expert's answer
2020-05-12T11:34:59-0400

Answer is 651 mmHg

Explanation:

Ptotal = P1 + P2 + ...

So P = 495+156=651